### v2

# A short proof of Isserlis’ theorem

Abstract.
We show that Isserlis’ theorem follows as a corollary to the invariant tensor theorem for isotropic tensors.
AMS Subject Classification (2020): Primary: 60C05, 15A72, Secondary: 41A58, 60H35.
1. Isserlis’s theorem
Isserlis’ theorem was first derived by Leon Isserlis in 1918 [7]. This result, also known as Wick’s probability theorem, permits to express arbitrary high-order moments of a given multivariate normal distribution ${\bf Y}$ in terms of the coefficients of its covariance matrix $\Sigma$.
Theorem 1 (Isserlis).
Let ${\bf Y}=(Y_{1},Y_{2},\ldots,Y_{n})^{T}\sim{\mathcal{N}}_{n}({\bf 0},\Sigma)$ be a multivariate normally distributed vector in ${\mathbb{R}}^{n}$ with mean zero. Then
(1) ${\mathsf{E}}(Y_{1}Y_{2}\cdots Y_{n})=\sum_{p\in\operatorname{PP}(n)}\prod_{(\ell,r)\in p}{\mathsf{E}}(Y_{\ell}Y_{r}).$
The sum is over all partitions of $n$ into disjoint pairs, denoted $\operatorname{PP}(n)$. For odd $n$, $\operatorname{PP}(n)=\emptyset$ is empty and (1) reduces to $E(Y_{1}Y_{2}\cdots Y_{n})=0$, and we recover that odd moments of normal distributions with mean zero always vanish. For even $n$ with $n=2k$, the number of elements of $\operatorname{PP}(n)$ is exactly #$\#\operatorname{PP}(n)=\frac{(2k)!}{2^{k}k!}.$ For $n=4$, we have #$\#\operatorname{PP}(4)=3$ with
$\operatorname{PP}(4)=\{\{(1,2),(3,4)\},\{(1,3),(2,4)\},\{(1,4),(2,3)\}\},$
and formula (1) becomes
${\mathsf{E}}(Y_{1}Y_{2}Y_{3}Y_{4})={\mathsf{E}}(Y_{1}Y_{2}){\mathsf{E}}(Y_{3}Y_{4})+{\mathsf{E}}(Y_{1}Y_{3}){\mathsf{E}}(Y_{2}Y_{4})+{\mathsf{E}}(Y_{1}Y_{4}){\mathsf{E}}(Y_{2}Y_{3}).$
This result was shown by Isserlis first for $n=4$ in [6] before he generalized the result for arbitrary integer $n$ in [7] with an elementary proof of (1) by induction.
2. Isserlis’ Theorem as a corollary of the invariant tensor theorem
Let $V,(\_,\_)$ be a Euclidean space (e.g. ${\mathbb{R}}^{d}$ with the standard inner product). A tensor $T\colon\otimes^{n}V\rightarrow{\mathbb{R}}$, where we denote $\otimes^{n}V=V\otimes V\otimes\cdots\otimes V$ ($n$ times), is called isotropic if $T(Q{\bf v}_{1},Q{\bf v}_{2},\ldots,Q{\bf v}_{n})=T({\bf v}_{1},\ldots,{\bf v}_{n})$ for all orthogonal matrices $Q\in O(V)$ and all vectors ${\bf v}_{i}\in V$. Note that the scalar product (_,_)__(\_,\_)( _ , _ ) itself is an example of an isotropic tensor $V\otimes V\rightarrow{\mathbb{R}}$.
We shall derive formula (1) as a consequence of the invariant tensor theorem for isotropic tensors [14, Sec. II.9][1, Appendix I][8, §33.2] stating that the linear space of isotropic tensors is spanned by certain elementary isotropic tensors with two inputs.
Theorem 2 (Invariant tensor theorem for isotropic tensors).
An isotropic tensor $T$ decomposes in a sum of products of pairwise inner products:
(2) $T({\bf v}_{1},{\bf v}_{2},\ldots,{\bf v}_{n})=\sum_{p\in\operatorname{PP}(n)}\alpha_{p}\prod_{(\ell,r)\in p}({\bf v}_{\ell},{\bf v}_{r})$
for some scalars $\alpha_{p}\in{\mathbb{R}}$ independent of ${\bf v}_{1},{\bf v}_{2},\ldots,{\bf v}_{n}$. If in addition, $T$ is symmetric, i.e. invariant under permutations of its inputs, then all $\alpha_{p}$ are equal, i.e. independent of $p$.
Proof.
The first part is a reformulation of [5, Th. 5.3.5]. The second part is obtained by contradiction. ∎
Again, for odd integers $n$, $\operatorname{PP}(n)=\emptyset$ and the right hand side in (2) vanishes.
We call a random variable ${\bf X}$ in ${\mathbb{R}}^{d}$ isotropic if its distribution is invariant by rotation, i.e., if for any $Q\in O(d)$ and any integrable function $f\colon{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$ we have ${\mathsf{E}}(f({\bf X}))={\mathsf{E}}(f(Q{\bf X}))$. For such a random variable, and all integer $k$, we define the quantities:
$c_{k}({\bf X}):=\frac{2^{k}k!}{(2k)!}\frac{{\mathsf{E}}(X^{2k})}{{\mathsf{E}}(X^{2})^{k}},$
where $X:={\bf a}^{T}{\bf X}$ for an arbitrary, nonzero, deterministic vector ${\bf a}\in{\mathbb{R}}^{d}$ (note that $c_{k}({\bf X})$ does not depend on the choice of ${\bf a}$ by the isotropy of ${\bf X}$). Note also that $c_{1}=1$.
We recall the following classical result on the covariance matrix of isotropic random variables.
Lemma 1.
Let ${\bf X}$ be an isotropic random variable in ${\mathbb{R}}^{d}$. Then ${\mathsf{E}}({\bf X}{\bf X}^{T})=\lambda I_{d}$ for some $\lambda\in{\mathbb{R}}$.
Proof.
The definition of isotropy with $f({\bf x})={\bf x}{\bf x}^{T}$ yields that the matrix $\Sigma={\mathsf{E}}({\bf X}{\bf X}^{T})$ commutes with all orthogonal matrices $Q\in O(d)$ a property satisfied only by matrices of the form $\lambda I_{d}$. Indeed, if a matrix $\Sigma$ commutes in particular with all orthogonal symmetries $S$, using that $\Sigma$ leaves invariant the eigenspace $\operatorname{Ker}(S+I_{d})$ of $S$, we obtain $\Sigma{\bf x}\in\operatorname{Vect}({\bf x})$ for all ${\bf x}\in{\mathbb{R}}^{d}$. This later property yields $\Sigma{\bf x}=\lambda{\bf x}$ for all ${\bf x}\in{\mathbb{R}}^{d}$ and for some $\lambda\in{\mathbb{R}}$ which is shown to be independent of ${\bf x}$ by contradiction and this concludes the proof. ∎
Theorem 3.
Let ${\bf X}$ be an isotropic random variable in ${\mathbb{R}}^{d}$. Then for any choice of $n$ deterministic vectors ${\bf a}_{1}$,…,${\bf a}_{n}\in{\mathbb{R}}^{d}$, we have
${\mathsf{E}}({\bf a}_{1}^{T}{\bf X}\cdots{\bf a}_{n}^{T}{\bf X})=\begin{cases}c_{k}({\bf X})\sum_{p\in\operatorname{PP}(n)}\prod_{(\ell,r)\in p}{\mathsf{E}}({\bf a}_{\ell}^{T}{\bf X}\,{\bf a}_{r}^{T}{\bf X}),&\text{if $n=2k$,}\\ 0&\text{if $n$ is odd.}\end{cases}$
Proof.
Let $A=({\bf a}_{1},{\bf a}_{2},\ldots,{\bf a}_{n})$ be the matrix containing the vectors ${\bf a}_{i}$ in its columns, and define $Y:=A^{T}{\bf X}$. As $X$ is isotropic, the tensor $T({\bf a}_{1},{\bf a}_{2},\ldots,{\bf a}_{n}):={\mathsf{E}}(Y_{1}Y_{2}\cdots Y_{n})={\mathsf{E}}({\bf a}_{1}^{T}{\bf X}\,{\bf a}_{2}^{T}{\bf X}\cdots{\bf a}_{n}^{T}{\bf X})$ is an isotropic and symmetric tensor, and hence of the form (2) for some real constant $\alpha_{p}=c$. We now compute ${\mathsf{E}}(Y_{\ell}Y_{r})={\mathsf{E}}({\bf a}_{\ell}^{T}{\bf X}\,{\bf a}_{r}^{T}{\bf X})={\mathsf{E}}({\bf a}_{\ell}^{T}{\bf X}{\bf X}^{T}{\bf a}_{r})={\bf a}_{\ell}^{T}{\mathsf{E}}({\bf X}{\bf X}^{T}){\bf a}_{r}=\lambda({\bf a}_{\ell},{\bf a}_{r})$ where we used Lemma 1. Finally, when $n=2k$, choosing ${\bf a}_{j}={\bf a}$, for an arbitrary vector ${\bf a}$ shows that $c=c_{k}({\bf X})$. ∎
Using that the moment generating function of the standard normal distribution is $\mathrm{e}^{t^{2}/2}$, one can show that ${\mathsf{E}}(X^{2k})=\frac{(2k)!}{2^{k}k!}=\#\operatorname{PP}(2k)$, from which we deduce for the case of multivariate normally distributed vectors with mean zero, $c_{k}({\bf X})=1$. This permits to deduce the proof of Isserlis’ Theorem 1.
3. Perspectives
Let us mention that the Isserlis theorem plays a fundamental role in the construction of exotic aromatic Butcher series, which are formal series index by certain graphs, introduced in [11, 12] for the numerical analysis of numerical integrators for ergodic stochastic dynamics. Such exotic aromatic Butcher series also yields new algebraic structures presented recently in [3, 4], generalizing aromatic Butcher series [2, 9] with a new kind of edge denoted lianas [11] and paring any two nodes, in the spirit of the notion of paring in Isserlis’ theorem. The construction of such lianas is based on the following result valid for an arbitrary tensor $T$ (without any isotropy or symmetry assumption on $T$), which can be proved as a straightforward consequence of the Isserlis theorem.
Theorem 4.
(see [11, Theorem 4.1]) Consider a normally distributed vector ${\bf Y}\sim{\mathcal{N}}_{d}({\bf 0},\Sigma)$ in ${\mathbb{R}}^{d}$ with mean zero and covariance matrix $\Sigma=AA^{T}$ with $A=({\bf a}_{1},\ldots,{\bf a}_{d})$. Then, for all tensor $T\colon\otimes^{n}{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$, we have
${\mathsf{E}}\big{(}T({\bf Y},{\bf Y},\ldots,{\bf Y})\big{)}=\sum_{p\in\operatorname{PP}(n)}\sum_{(k_{1},k_{2},\ldots,k_{n})\in K_{d}(p)}T({\bf a}_{k_{1}},\ldots,{\bf a}_{k_{n}})$
where for each pairing $p\in PP(n)$, we define the set of paired indices
$K_{d}(p):=\{(k_{1},\ldots,k_{n})\in\{1,\ldots,d\}^{n}\ ;\ k_{\ell}=k_{r}\mbox{for all }(\ell,r)\in p\}.$
Note that for all $p\in\operatorname{PP}(2k)$, the set $K_{d}(p)$ has $d^{k}$ elements. For instance, for $n=4$,
${\mathsf{E}}\big{(}T({\bf Y},{\bf Y},{\bf Y},{\bf Y})\big{)}=\sum_{i,j=1}^{d}\big{(}T({\bf a}_{i},{\bf a}_{i},{\bf a}_{j},{\bf a}_{j})+T({\bf a}_{i},{\bf a}_{j},{\bf a}_{i},{\bf a}_{j})+T({\bf a}_{i},{\bf a}_{j},{\bf a}_{j},{\bf a}_{i})\big{)}.$
We believe that this short proof of Isserlis’ theorem using the invariant tensor theorem for isotropic tensors could give new insight on further studies of equivariance characterization properties for Butcher series [13] and generalisations [10].
Acknowledgements: We thank Arieh Iserles for enlightening us on the genealogy of Moses Isserles. H. M.-K. is supported by the Research Council of Norway through project 302831 Computational Dynamics and Stochastics on Manifolds (CODYSMA). G.V. was partially supported by the Swiss National Science Foundation, projects No 200020_214819 and No. 200020_192129.